OpenStudy (anonymous):

. Simplify the following expression, and rewrite it in an equivalent form with positive exponents. -30x^4y^-4 / 120^2y^-4

OpenStudy (anonymous):

But with what you have given me then your answer will be -x^4/480

OpenStudy (anonymous):

no it is not supposed to be an x==> sorry i didnt see your response.

OpenStudy (anonymous):

so how ould the problem be rewritten?

OpenStudy (anonymous):

the exponents are positive...

OpenStudy (anonymous):

the complete answer is rewritten and simplified how ? i am sorry.

OpenStudy (anonymous):

oh sorry...

OpenStudy (anonymous):

\[\frac{\frac{-30x^4}{y^4}}{\frac{120^2}{y^4}}\]

OpenStudy (anonymous):

That is the expression rewritten with positive exponents but the answer I gave you is it simplified do you understand how to get to the answer?

OpenStudy (anonymous):

the y^4's cancel each other

OpenStudy (anonymous):

or how it was originally \[\frac{-30x^4y^(-4+4))}{480}\]

OpenStudy (anonymous):

not 480 14400

OpenStudy (anonymous):

so the last one is the answer? in the divison form? then its -30x^4y -(-4 + 4 / 1400

OpenStudy (anonymous):

14400

OpenStudy (anonymous):

no... -x^4/480 is the final simplified answer with positive exponents... the x is raised to a positive power

OpenStudy (anonymous):

ok so the answer would be written out how? I think I am way off base :(

OpenStudy (anonymous):

Is this in a text book? If so, can you check the solution if not... the answer to my knowledge is written \[-\frac{x^4}{480}\]

OpenStudy (anonymous):

it was 3 text book questions that were emailed to me and I was way off. thanks

OpenStudy (anonymous):

You're welcome! All the best :).

OpenStudy (anonymous):

thanks appreciate it.